How many atoms are in a 51 g sample of gold?


1
Expert's answer
2021-03-08T05:58:23-0500

mass of sample of gold=591gm

molar mass of gold=197g/mol

"mole= \tfrac{mass}{molar mass}\\

mole=\frac{591}{197}\\

mole= 3\\

Total number of atoms= mole* 6*10^{23}\\

Atoms=3*6*10^{23}\\

Total number of atoms= 18*10^{23}"